Subtract the indicated bases.
2102(three)-112(three)

Answer by Alan3354(69443)

Subtract the indicated bases.
2102(three)-112(three)
-----------------
2102
-112
----------
1220
=========
2102 = 2*27 + 9 + 2 = 65 decimal
112 = 9 + 3 + 2 = 14 decimal
--
1220 = 27 + 18 + 6 = 51 decimal
65 - 14 = 51